Assume that when adults with smartphones are randomly​ selected, 61​% use them in meetings or classes. If 7 adult smartphone users are randomly​ selected, find the probability that exactly 3 of them use their smartphones in meetings or classes.


1
Expert's answer
2021-02-28T08:29:03-0500

p = 61 % = 0.61

n = 7

"P(X=x) = \frac{n!}{x!(n-x)!} \times p^x \times (1-p)^{n-x} \\

p(X=3) = \frac{7!}{3!(7-3)!} \times 0.61^3 \times (1-0.61)^{7-3} \\

= 42 \times 0.61^3 \times 0.39^4 \\

= 0.160"